# FreeTacMan: Robot-free Visuo-Tactile Data Collection System for Contact-rich Manipulation

### v2

@@ Abstract
Enabling robots with contact-rich manipulation remains a pivotal challenge in robot learning, which is substantially hindered by the data collection gap, including its inefficiency and limited sensor setup. While prior work has explored handheld paradigms, their rod-based mechanical structures remain rigid and unintuitive, providing limited tactile feedback and posing challenges for human operators. Motivated by the dexterity and force feedback of human motion, we propose FreeTacMan, a human-centric and robot-free data collection system for accurate and efficient robot manipulation. Concretely, we design a wearable gripper with dual visuo-tactile sensors for data collection, which can be worn by human fingers for intuitive control. A high-precision optical tracking system is introduced to capture end-effector poses while synchronizing visual and tactile feedback simultaneously. We leverage FreeTacMan to collect a large-scale multimodal dataset, comprising over 3000k paired visual-tactile images with end-effector poses, 10k demonstration trajectories across 50 diverse contact-rich manipulation tasks. FreeTacMan achieves multiple improvements in data collection performance compared to prior works, and enables effective policy learning for contact-rich manipulation tasks with self-collected dataset. The full suite of hardware specifications and the dataset will be released to facilitate reproducibility and support research in visuo-tactile manipulation.
@@ I INTRODUCTION
The vision-based imitation learning has shown strong potential in robot manipulation tasks [5, 6], benefiting from the growing large-scale demonstration datasets [7, 8]. In tactile domain, visuo-tactile sensors offer high-resolution, sensitive, and easily integrable multimodal signals, making them particularly well-suited for existing policy learning pipelines. However, the lack of large-scale, high-quality tactile datasets and compatible sensing hardware has prevented similar advances. To accelerate research in this domain, two prerequisites would be addressed: 1) a visuo-tactile data collection system that provides real-time tactile feedback and enables rapid redeployment, and 2) a large-scale, high-precision visuo-tactile dataset covering diverse contact-rich manipulation tasks.
Early efforts concentrated on collecting data using sensors mounted on robots [9, 10, 11], which limit the flexibility for scaling visuo-tactile data. Systems that rely on motion-capture with AR/VR visualization [12, 13] or primary-replica teleoperation rigs [14] capture accurate camera views and trajectories. They conversely offer no direct, real-time tactile signals, and impose fixed robot setup with complex calibration or high latency. Handheld data collection paradigms [15, 16] release the human operators from robot embodiment. However, these methods typically rely on SLAM and IMU fusion for localization, which introduces significant positioning errors that are particularly detrimental to tactile perception. Moreover, they transmit tactile cues through long mechanical linkages, hindering direct tactile feedback and precise measurement of instantaneous tactile signals.
The inefficiency and lack of real-time tactile feedback in current data collection setups compromise data quality, limit scalability, and pose a risk of sensor damage during operation. Consequently, existing visuo-tactile datasets[17, 18, 19] are largely confined to specific perception tasks or lack the diversity and precision required for generalizable visuo-tactile policy learning across diverse manipulation scenarios.
@@
TABLE I: Comparison with existing data collection systems. Our in-situ design delivers direct and high-precision tactile feedback to operators. To evaluate tactile feedback fidelity of handheld methods quantitatively, we measure the number of mechanical transmission “link” inside the handheld gripper from human hand to the grasped object.
@@
Teleop:
VR/AR
ARCap [12] VR Controller
@@
Teleop:
Primary–
Replica
ALOHA [14] Puppet Arm
@@
Handheld
UMI [15]
Trigger
Contact (4 links)
FastUMI [16]
@@ II Related Work
Recent robot learning has been greatly advanced through imitating expert demonstrations. Popular approaches [26, 27, 28] leverage large-scale visuomotor datasets (pairs of RGBD sensors and actions) to train end-to-end policies that generalize across objects and scenes. To gather data suitable for imitation learning, prior works have relied on different interfaces to teleoperate real robots, including motion capture with AR or VR visualization [12, 8, 13], 3D space mouse [27], primary-replica system [14, 24, 25], and wearable devices [29]. However, these setups are either expensive, operationally complex, or suffer from limited precision. Handheld data collection paradigms [15] offer greater flexibility for various robot embodiments, but their trigger-based grippers and multi-link transmission designs introduce backlash between links, which blurs tactile cues and leads to inaccurate relay of gripper contact to the operator. In contrast, FreeTacMan implements an in-situ data collection system that enables operators to feel gripper contact directly in real time. The comparison of control methods and tactile feedback between ours and prior work is depicted in Table˜I.
@@
Most existing visuo-tactile datasets focus on perception tasks such as 6DoF pose tracking[30], cross-modal generation[17], representation learning[19], garment-feature spatially-aligned perception[31], and tactile-semantic description[32], with limited coverage of the full manipulation process. The ObjectFolder benchmark [33] includes 4 visuo-tactile manipulation tasks, exhibiting limited scale and diversity. Moreover, most tactile manipulation datasets are collected with tactile array sensors based on piezoresistive or Hall-effect principles [34, 35, 36, 37]. Although policies trained on these dataset show promising results, the intrinsic limitations of these sensors, such as low spatial resolution, crosstalk, complex fabrication, and environment interference, can constrain policy performance and reduce deployment efficiently in real-world settings. Taken together, these limitations expose an urgent demand for a visuo-tactile dataset that offers rich, high-resolution, and scalable data spanning the entire manipulation process.
@@ III-A Hardware Design
Multimodal sensing. As shown in Figure˜2(a), we employ visuo-tactile sensors and a wrist-mounted camera to capture tactile and visual information. End-effector poses are tracked using a high-precision motion capture system, achieving sub-millimeter accuracy—an essential factor for contact-rich tasks, where even minute pose errors can result in disproportionately large deviations in tactile feedback.
@@
Modular architecture. FreeTacMan system is built as three plug-and-play modules, each optimized for rapid setup and cross-embodiment compatibility: a sensor perception module for tactile data collection, a universal gripper interface (Fig.˜2(c)) for robot compatibility , and a camera mounting scaffold to ensure aligned visual feedback from wrist camera. The sensor is based on the McTac design [38] and features enhanced modularity and consistency. The improved universal mechanical interface allows the same sensor unit to be used interchangeably on a data collection setup or a robotic arm. Automated fabrication ensures consistent quality. Customized interfaces are provided for different robot arms, including a 6-DOF Piper arm for low-load tasks and 7-DOF Franka arm for high-precision, heavy-load applications. For 3D models of end-effectors and their integration on different robot arms (e.g., Piper, Franka), please visit our demo page.
@@ III-B Human-to-Robot Data Transfer
To facilitate human-to-robot skill transfer, a high-precision motion capture system [40] is used for 6D pose tracking of the interface at 300 Hz, achieving a mean error of 0.118 mm. Five retro-reflective markers are mounted on the interfaces, with three positioned on the top plate to measure pose and two on the grippers to capture relative displacement. The coordinates of markers are first transformed from the world coordinate frame to the robot base coordinate frame. A local coordinate frame, aligned with the robot URDF, is established at the gripper’s Tool Center Point (TCP) using three top-plate markers, allowing derivation of end-effector pose while ensuring consistency with the robot kinematic model. We downsample the tracking data to synchronize with RGB images. To this end, each frame contains: the wrist camera RGB image, two visuo-tactile images, the end-effector pose of the in-situ gripper in the world coordinate frame and gripper width. Each trajectory consists of a sequence of such embodiment-agnostic data synchronized at 30 Hz.
@@ IV Experiments
Q3. How dose the tactile encoder pretrained on self-collected visuo-tactile data improve policy learning?
@@ IV-A Dataset
Figure 4: The FreeTacMan dataset. (a) Representative examples illustrating the diversity in both task complexity and tactile context. (b) The dataset covers 50 tasks and features a large-scale collection of data, including more than 10k trajectories which contains over 3000k visuo-tactile pairs. (c) The dataset enables diverse fundamental tactile capabilities.
Enabled by the efficient, precise, and fidelity tactile data collection system, we curate a diverse dataset of manipulation tasks spanning vision, touch, and proprioception modalities, as illustrated in Fig.˜4. The dataset spans 50 tasks, comprising more than 10k manipulation trajectories which contains over 3000k visuo-tactile image pairs. Collected with high-consistent and high-resolution visuo-tactile sensors, it enables large-scale tactile pretraining and multimodal policy learning. Furthermore, the dataset covers diverse fundamental tactile capabilities, as shown in Figure˜4(c). To ensure dataset quality, every trajectory is replayed in our validation process, providing a robust filter that guarantees the reliability of the data for downstream robot learning.
@@ IV-B Experimental Setup
Figure 5: Human demonstrations and policy rollouts. The top row shows goal trajectories, the middle row demonstrates successful rollouts with tactile feedback, and the bottom row showcases typical failure modes when without tactile input.
To answer the questions above, we evaluate the effectiveness of FreeTacMan system and the quality of the collected dataset through a diverse set of contact-rich manipulation tasks, shown in Figure˜5. 1) Fragile cup. The robot grasps a fragile cup without causing damage and places it stably on a tray. 2) Calligraphy. The robot traces the digit "5" with a calligraphy brush. 3) Texture classification. The robot grasps and identifies one of two cylindrical objects with distinct textures and sorts it into the correct bin. 4) Stamp press. The robot presses a stamp onto paper to produce a clear imprint. 5) USB plug. The robot needs to securely insert a pre-grasped USB plug into a socket (error <1mm<1mm). These tasks collectively represent the four core tactile capabilities of our dataset: force control (fragile cup), hybrid force-position control (USB plug, stamp press), high-acuity tactile perception (texture classification) and dynamic response (calligraphy).
@@ IV-C User Study on Data Collection System
Method
Fragile
Cup
Calligraphy
Texture
Cls.
Stamp
Press
USB
Plug
Avg.
ACT [26]
(Vision-only)
35 20 20 30 0 21
Ours (+ Tactile
w/o Pretrain)
75 70 55 65 10 55
Ours
(+ Pretrain)
80 90 85 80 20 71
@@ IV-D Validation on Imitation Learning
Integration of tactile feedback results in significant improvement on policy performance (Q2). As illustrated in Table˜II, the vision-only baseline achieves low performance across all tasks, with an average success rate of 21%. Without tactile feedback, the robot struggles in tasks requiring fine tactile perception, such as USB plug, calligraphy, and texture classification. When tactile feedback is incorporated naively, i.e., without pretraining, performance improves substantially, with the success rate increasing to 55%. Significant gains are observed in tasks such as fragile cup manipulation (75%) and calligraphy (70%). This highlights the utility of tactile sensing in contact-rich tasks where visual cues alone are insufficient to distinguish subtle features. However, the performance in tasks like USB plug (10%) and texture classification (55%) still lags behind, indicating room for further refinement.
Tactile inputs enables excellent performance even on unseen objects (Q2). In the texture classification task, we select a texture object with an out-of-distribution color (red) for testing. As shown in Table˜III, tactile input achieves 55% accuracy on unseen objects, matching performance on training objects and significantly exceeding the 15% vision-only baseline. Pretraining further increases accuracy to 70%, demonstrating robust generalization.
@@
Temporal-aware pretraining improves the performance by aligning visual and tactile embeddings (Q3). Incorporating both time-aligned and time-adjacent visual-tactile pairs in CLIP pretraining leads to a notable performance boost, with the average success rate increasing to 71%. As shown in the last line in Table˜II, the improvement comes from tasks that demand fine-grained control and the ability to track contact dynamics, such as calligraphy (90%) and stamp (80%). Meaningful gain is witnessed in USB plug (20%). These results show that our approach aligns tactile and visual embeddings while incorporating spatial and temporal context.
@@
Figure 8: Ablation across training episodes Our pretrained policy, even with small data, outperforms vision-only policy trained with large amounts of data. Moreover, it approaches the performance of a large-data tactile-only policy.
Pretraining on the FreeTacMan dataset provides robustness against variations among visuo-tactile sensors (Q3). While our main experiments are conducted in an in-domain setting, where training and evaluation use different sensor batches with high consistency, we also evaluate a more challenging out-of-domain scenario by introducing substantial variations across visuo-tactile sensors. As shown in Table˜IV, pretrained model generalizes well across two distinct sensors, which differ in marker angle (0° vs. 45°) and RGB lighting environment (side vs. bottom). The in-domain and out-of-domain performance remain close on the texture classification task (0.85 vs. 0.75) and identical on the fragile cup task (both 0.8). In contrast, the non-pretrained counterpart performs poorly, particularly in phases requiring fine-grained tactile perception: a drop from 80% to 40% in the tactile-based classification phase, and a catastrophic drop from 70% to 10% in the landing phase (which relies on detecting contact with table before releasing). These results validate the effectiveness of tactile pretraining for cross-sensor generalization.

### v3

@@ Abstract
Enabling robots with contact-rich manipulation remains a pivotal challenge in robot learning, which is substantially hindered by the data collection gap, including its inefficiency and limited sensor setup. While prior work has explored handheld paradigms, their rod-based mechanical structures remain rigid and unintuitive, providing limited tactile feedback and posing challenges for operators. Motivated by the dexterity and force feedback of human motion, we propose FreeTacMan, a human-centric and robot-free data collection system for accurate and efficient robot manipulation. Concretely, we design a wearable gripper with visuo-tactile sensors for data collection, which can be worn by human fingers for intuitive control. A high-precision optical tracking system is introduced to capture end-effector poses while synchronizing visual and tactile feedback simultaneously. We leverage FreeTacManto collect a large-scale multimodal dataset, comprising over 3000k paired visuo–tactile images with end-effector poses, 10k demonstration trajectories across 50 diverse contact-rich manipulation tasks. FreeTacManachieves multiple improvements in data collection performance over prior works and enables effective policy learning from self-collected datasets. By open-sourcing the hardware and the dataset, we aim to facilitate reproducibility and support research in visuo-tactile manipulation.
@@ I INTRODUCTION
The vision-based imitation learning has shown strong potential in robot manipulation tasks [5, 6, 7, 8], benefiting from the growing large-scale demonstration datasets [9, 10]. In the tactile domain, visuo-tactile sensors offer high-resolution, sensitive, and easily integrable multimodal signals, making them particularly well-suited for existing policy learning pipelines. However, the lack of large-scale, high-quality tactile datasets and compatible sensing hardware has prevented similar advances. To accelerate research in this domain, two prerequisites would be addressed: 1) a visuo-tactile data collection system that provides real-time tactile feedback and enables rapid redeployment, and 2) a large-scale, high-precision visuo-tactile dataset covering diverse contact-rich manipulation tasks [11].
Early efforts concentrated on collecting data using sensors mounted on robots [12, 13], which limit the flexibility for scaling visuo-tactile data. Systems that rely on motion-capture with AR/VR visualization [14] or primary-replica teleoperation rigs [15] capture accurate camera views and trajectories. They conversely offer no direct, real-time tactile signals, and impose a fixed robot setup with complex calibration or high latency. Handheld data collection paradigms [16, 17] release the human operators from robot embodiment. However, these methods typically rely on SLAM and IMU fusion for localization, which introduces significant positioning errors that are particularly detrimental to tactile perception. Moreover, they transmit tactile cues through long mechanical linkages, hindering direct tactile feedback and precise measurement of instantaneous tactile signals.
The inefficiency and lack of real-time tactile feedback in current data collection setups compromise data quality, limit scalability, and pose a risk of sensor damage during operation. Consequently, existing visuo-tactile datasets[18, 19] are largely confined to specific perception tasks or lack the diversity and precision required for generalizable visuo-tactile policy learning across diverse manipulation scenarios.
@@
TABLE I: Comparison with existing data collection systems. Our in-situ design delivers direct and high-precision tactile feedback to operators. To evaluate tactile feedback fidelity of handheld methods quantitatively, we measure the number of mechanical transmission “link” inside the handheld gripper from the human hand to the grasped object.
@@
Teleop: VR/AR ARCap [22] VR Controller
@@
Teleop: Primary– Replica ALOHA [15] Puppet Arm
@@
Handheld UMI [16], FastUMI [17] Trigger Contact (4 links)
ViTaMIn [27]
@@ II Related Work
Recent robot learning has been greatly advanced through imitating expert demonstrations [28]. Popular approaches [29, 30] leverage large-scale visuomotor datasets (pairs of RGBD sensors and actions) to train end-to-end policies that generalize across objects and scenes. To gather data suitable for imitation learning, prior works have relied on different interfaces to teleoperate real robots, including motion capture with AR or VR visualization [22, 10, 14], 3D space mouse [30], primary-replica system [15, 25, 26], and wearable devices [31]. However, these setups are either expensive, operationally complex, or suffer from limited precision. Handheld data collection paradigms [16] offer greater flexibility for various robot embodiments, but their trigger-based grippers and multi-link transmission designs introduce backlash between links, which blurs tactile cues and leads to inaccurate relay of gripper contact to the operator. In contrast, FreeTacManimplements an in-situ data collection system that enables operators to feel gripper contact directly in real time. The comparison of control methods and tactile feedback between ours and prior work is depicted in Table˜I.
@@
Most existing visuo-tactile datasets focus on perception tasks such as 6DoF pose tracking [32], cross-modal generation [18], representation learning [19], garment-feature spatially-aligned perception [33], and tactile-semantic description [34], with limited coverage of the full manipulation process. The ObjectFolder benchmark [35] includes 4 visuo-tactile manipulation tasks, exhibiting limited scale and diversity. Moreover, most tactile manipulation datasets are collected with tactile array sensors based on piezoresistive or Hall-effect principles [13, 36, 37, 38]. Although policies trained on these datasets show promising results, the intrinsic limitations of these sensors, such as low spatial resolution, crosstalk, complex fabrication, and environment interference, can constrain policy performance and reduce deployment efficiency in real-world settings. Taken together, these limitations expose an urgent demand for a visuo-tactile dataset that offers rich, high-resolution, and scalable data spanning the entire manipulation process.
@@ III-A Hardware Design
Multimodal sensing. As shown in Fig.˜2(a), we employ visuo-tactile sensors and a wrist-mounted camera to capture tactile and visual information. End-effector poses are tracked using a high-precision motion capture system, achieving sub-millimeter accuracy—an essential factor for contact-rich tasks, where even minute pose errors can result in disproportionately large deviations in tactile feedback.
@@
Modular architecture. FreeTacMansystem is built as three plug-and-play modules, each optimized for rapid setup and cross-embodiment compatibility: a sensor perception module for tactile data collection, a universal gripper interface (Fig.˜2(c)) for robot compatibility, and a camera mounting scaffold to ensure aligned visual feedback from wrist camera. The sensor is based on the McTac design [39] and features enhanced modularity and consistency. The improved universal mechanical interface allows the same sensor unit to be used interchangeably on a data collection setup or a robotic arm. Automated fabrication ensures consistent quality. Customized interfaces are provided for different robot arms, including a 6-DOF Piper arm for low-load tasks and a 7-DOF Franka arm for high-precision, heavy-load applications. For 3D models of end-effectors and their integration on different robot arms (e.g., Piper, Franka), please visit our demo page.
@@ III-B Human-to-Robot Data Transfer
To facilitate human-to-robot skill transfer, a high-precision NOKOV Motion Capture System is used for 6D pose tracking of the interface at 240 Hz, achieving sub-millimeter accuracy. Five retro-reflective markers are mounted on the interfaces, with three positioned on the top plate to measure pose and two on the grippers to capture relative displacement. The coordinates of markers are first transformed from the world coordinate frame to the robot base coordinate frame. A local coordinate frame, aligned with the robot URDF, is established at the gripper’s Tool Center Point (TCP) using three top-plate markers, allowing derivation of end-effector pose while ensuring consistency with the robot kinematic model. We downsample the tracking data to synchronize with RGB images. To this end, each frame contains: the wrist camera RGB image, two visuo-tactile images, the end-effector pose of the in-situ gripper in the world coordinate frame, and gripper width. Each trajectory consists of a sequence of such embodiment-agnostic data synchronized at 30 Hz.
@@ IV Experiments
Q3. How does the tactile encoder pretrained on self-collected visuo-tactile data improve policy learning?
@@ IV-A Dataset
Figure 4: The FreeTacMandataset. (a) Representative examples illustrating the diversity in both task complexity and tactile context. (b) The dataset covers 50 tasks and features a large-scale collection of data, including more than 10k trajectories and over 3000k visuo-tactile pairs. (c) The dataset enables diverse fundamental tactile capabilities.
Enabled by the efficient, precise, and faithful tactile data collection system, we curate a diverse dataset spanning vision, touch, and proprioception modalities, as illustrated in Fig.˜4. The dataset spans 50 tasks, comprising more than 10k trajectories and over 3000k visuo-tactile image pairs. Collected with high-consistent and high-resolution visuo-tactile sensors, it enables large-scale tactile pretraining and multimodal policy learning. Furthermore, the dataset covers diverse fundamental tactile capabilities, as shown in Fig.˜4(c). To ensure dataset quality, every trajectory is replayed in our validation process, providing a robust filter that guarantees the reliability of the data for downstream robot learning. The dataset, along with its structure, data format, and license, is available at our dataset page.
@@ IV-B Experimental Setup
Figure 5: Human demonstrations and policy rollouts. The top row shows goal trajectories, the middle row demonstrates successful rollouts with tactile feedback, and the bottom row showcases typical failure modes without tactile input.
To answer the questions above, we evaluate the effectiveness of FreeTacMansystem and the quality of the collected dataset through a diverse set of contact-rich manipulation tasks, shown in Fig.˜5. 1) Fragile cup. The robot grasps a fragile cup without causing damage and places it stably on a tray. 2) Calligraphy. The robot traces the digit "5" with a calligraphy brush. 3) Texture classification. The robot grasps and identifies one of two cylindrical objects with distinct textures and sorts it into the correct bin. 4) Stamp press. The robot presses a stamp onto paper to produce a clear imprint. 5) USB plug. The robot needs to securely insert a pre-grasped USB plug into a socket (error <1mm<1mm). These tasks collectively represent the four core tactile capabilities of our dataset: force control (fragile cup), hybrid force-position control (USB plug, stamp press), high-acuity tactile perception (texture classification), and dynamic response (calligraphy).
@@ IV-C User Study on Data Collection System
Method Fragile Cup Calligraphy Texture Cls. Stamp Press USB Plug Avg.
ACT [29] (Vision-only) 35 20 20 30 0 21
Ours (+ Tactile w/o Pretrain) 75 70 55 65 10 55
Ours (+ Pretrain) 80 90 85 80 20 71
@@ IV-D Validation on Imitation Learning
Integration of tactile feedback results in significant improvements on policy performance (Q2). As illustrated in Table˜II, the vision-only baseline achieves low performance across all tasks, with an average success rate of 21%. Without tactile feedback, the robot struggles in tasks requiring fine tactile perception, such as USB plug, calligraphy, and texture classification. When tactile feedback is incorporated naively, i.e., without pretraining, performance improves substantially, with the success rate increasing to 55%. Significant gains are observed in tasks such as fragile cup manipulation (75%) and calligraphy (70%). This highlights the utility of tactile sensing in contact-rich tasks where visual cues alone are insufficient to distinguish subtle features. However, the performance in tasks like USB plug (10%) and texture classification (55%) still lags behind, indicating room for further refinement.
Tactile inputs enable excellent performance even on unseen objects (Q2). In the texture classification task, we select a texture object with an out-of-distribution color (red) for testing. As shown in Table˜III, tactile input achieves 55% accuracy on unseen objects, matching performance on training objects and significantly exceeding the 15% vision-only baseline. Pretraining further increases accuracy to 70%, demonstrating robust generalization.
@@
Temporal-aware pretraining improves the performance by aligning visual and tactile embeddings (Q3). Incorporating both time-aligned and time-adjacent pairs in CLIP pretraining boosts the average success rate to 71%. As shown in the last line in Table˜II, the improvement comes from tasks that demand fine-grained control and the ability to track contact dynamics, such as calligraphy (90%) and stamp (80%). Meaningful gain is witnessed in USB plug (20%). Nevertheless, this 20% success rate highlights the challenge of high-precision tasks, where sub-millimeter deviations lead to failure. This is attributed to limited inverse kinematics accuracy and insufficient modeling of insertion dynamics. Future work will thus explore higher-precision arms and reinforcement learning to better learn contact-rich policies. Overall, these results confirm that our approach aligns tactile and visual embeddings with spatial-temporal context.
@@
Figure 8: Ablation across training episodes. Our pretrained policy, even with small data, outperforms the vision-only policy trained with large amounts of data. Moreover, it approaches the performance of a large-data tactile-only policy.
@@
Pretraining on FreeTacMandataset provides robustness against variations among visuo-tactile sensors (Q3). While our main experiments are conducted in an in-domain setting, where training and evaluation use different sensor batches with high consistency, we also evaluate a more challenging out-of-domain scenario by introducing substantial variations across visuo-tactile sensors. As shown in Table˜IV, the pretrained model generalizes well across two distinct sensors, which differ in marker angle (0° vs. 45°) and RGB lighting environment (side vs. bottom). The in-domain and out-of-domain performances remain close on the texture classification task (0.85 vs. 0.75) and identical on the fragile cup task (both 0.8). In contrast, the non-pretrained counterpart performs poorly, particularly in phases requiring fine-grained tactile perception: a drop from 80% to 40% in the tactile-based classification phase, and a catastrophic drop from 70% to 10% in the landing phase (which relies on detecting contact with the table before releasing). These results validate the effectiveness of tactile pretraining for cross-sensor generalization.